OpenStudy (anonymous):

lim - > 2 1/2 - 1/x divided by 2-x

OpenStudy (anonymous):

I got up to x-2/2x divided by 2-x , not sure what to do after

geerky42 (geerky42):

\[\lim_{x\rightarrow2}\dfrac{\dfrac{x-2}{2x}}{2-x} = \lim_{x\rightarrow2}\dfrac{x-2}{2x(2-x)}\] Right? Do you know what to do next?

geerky42 (geerky42):

HINT: \(x-2 = -(2-x)\)